Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

its an easy question

Expert replies
by sana.noor » Mon Sep 23, 2013 10:35 pm
i am trying to solve this question in two different ways, one way is easy and i get right answer but the other way is lengthy but i am not getting the right answer, am i doing it right? here is the question and explanation
there are 100 balls in a jar. 40 are black and 60 are white. what is the probability of getting at least one black in three consecutive picks with replace of each ball.
a) .216
b).316
c).5
d).684
e).784
OA is 0.784

one way: 1-(getting no black ball) = 1- (3/5)^3= 0.784 answer

2nd way: 1) getting one black and 2 white= 18/125
2)getting two black and one white = 12/125
3) getting 3 black = 8/125

adding all three cases = 38/125 =.304 and thats the wrong answer, what i am doing wrong here
Work hard in Silence, Let Success make the noise.

If you found my Post really helpful, then don't forget to click the Thank/follow me button. :)
Join the discussion
Source: — Problem Solving |

by vipulgoyal » Mon Sep 23, 2013 10:45 pm
just use idiot, i mean 5!/2!

multiply eash fragment by 3,3,1
54/125 + 36/125 + 8/125 = .784
Join the discussion

by [email protected] » Mon Sep 23, 2013 11:32 pm
Hi sana.noor,

Since the question asks for the odds of AT LEAST ONE black in 3 picks, the "easy" way to solve this problem is to calculate what you DON'T WANT and then subtract from 1:

Odds of a Black ball = 2/5
Odds of a White ball = 3/5

1 - (NO BLACK BALLS) = 1 - (3/5)(3/5)(3/5) = 1 - 27/125 = 98/125 = .784

As a decimal, it would be 1 - (.6)(.6)(.6) = 1 - .216 = .784

The LONG WAY is to calculate ALL possibilities and add them up:

1 Black and 2 White (any 1 could be black so x3) = (.4)(.6)(.6) x3 = .144x3 = .432
2 Black and 1 White (any 1 could be white so x3) = (.4)(.4)(.6) x3 = .096 = .288
3 Blacks and 0 White ---------> = (.4)(.4)(.4) = .064

.432 + .288 + .064 = .784

GMAT assassins aren't born, they're made,
Rich
Last edited by [email protected] on Tue Sep 24, 2013 6:21 pm, edited 1 time in total.
Contact Rich at [email protected]
Image
Join the discussion

by theCodeToGMAT » Mon Sep 23, 2013 11:36 pm
Sana, you just took the Patern as 1B2W without considering the different possible patterns formed by them.
For example:
BWW
WBW
WWB
So, you need to multiply Step 1 & 2 for Second Way by 3 to get the answer

My Steps:

3(4 x 6 x 6)/1000 + 3(4 x 4 x 6)/1000 + (4 x 4 x 4)/1000
= 784/1000 = .784
Last edited by theCodeToGMAT on Tue Sep 24, 2013 1:25 am, edited 1 time in total.
R A H U L
Join the discussion

by ganeshrkamath » Tue Sep 24, 2013 12:34 am
sana.noor wrote:i am trying to solve this question in two different ways, one way is easy and i get right answer but the other way is lengthy but i am not getting the right answer, am i doing it right? here is the question and explanation
there are 100 balls in a jar. 40 are black and 60 are white. what is the probability of getting at least one black in three consecutive picks with replace of each ball.
a) .216
b).316
c).5
d).684
e).784
OA is 0.784

one way: 1-(getting no black ball) = 1- (3/5)^3= 0.784 answer

2nd way: 1) getting one black and 2 white= 18/125
2)getting two black and one white = 12/125
3) getting 3 black = 8/125

adding all three cases = 38/125 =.304 and thats the wrong answer, what i am doing wrong here
Probability of getting at least 1 black ball = 1 - probability of getting no black balls
= 1 - (6/10)(6/10)(6/10)
= 1 - (0.6 * 0.6 * 0.6)
= 1 - 0.216
= 0.784

Choose e

The other way:
p(one black ball) = (4/10)(6/10)(6/10) * 3C1____________________3C1 because the order matters
= 0.144*3
= 0.432
p(2 black balls) = (4/10)(4/10)(6/10) * 3C2
= 0.096*3
= 0.288
p(3 black balls) = (4/10)(4/10)(4/10)
= 0.064

So p(at least one black ball) = 0.432 + 0.288 + 0.064 = 0.784

Choose e

Cheers
Every job is a self-portrait of the person who did it. Autograph your work with excellence.

Kelley School of Business (Class of 2016)
GMAT Score: 750 V40 Q51 AWA 5 IR 8
https://www.beatthegmat.com/first-attemp ... tml#688494
Join the discussion

by GMATGuruNY » Tue Sep 24, 2013 12:57 am
sana.noor wrote:i am trying to solve this question in two different ways, one way is easy and i get right answer but the other way is lengthy but i am not getting the right answer, am i doing it right? here is the question and explanation
there are 100 balls in a jar. 40 are black and 60 are white. what is the probability of getting at least one black in three consecutive picks with replace of each ball.
a) .216
b).316
c).5
d).684
e).784
OA is 0.784

one way: 1-(getting no black ball) = 1- (3/5)^3= 0.784 answer

1) getting one black and 2 white= 18/125
2) getting two black and one white = 12/125

3) getting 3 black = 8/125

adding all three cases = 38/125 =.304 and thats the wrong answer, what i am doing wrong here
Case 1: exactly 1 black marble
You have correctly determined P(BWW):
The probability that the FIRST marble is black, while the second and third marbles are white.
But this is not the ONLY way to get exactly 1 black marble.
We must account for ALL of the ways to get exactly 1 black marble.
Since the 1 black marble could be selected on any of the 3 picks -- BWW, WBW, WWB -- we multiply by 3:
3 * 18/125 = 54/125.

Case 2: exactly 2 black marbles, implying exactly 1 white marble
You have correctly determined P(BBW):
The probability that the FIRST AND SECOND MARBLES are black, while the third marble is white.
But this is not the ONLY way to get exactly 2 black marbles.
We must account for ALL of the ways to get exactly 2 black marbles.
Since the 1 white marble could be selected on any of the 3 picks -- WBB, BWB, BBW -- we multiply by 3:
3 * 12/125 = 36/125.

Case 3: all 3 marbles are black
You have correctly determined P(BBB) = 8/125.
Since there is only one way to get exactly 3 black marbles -- BBB -- we don't need to multiply here.

Adding the fractions, we get:

54/125 + 36/125 + 8/125 = 98/125 = .784.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by sana.noor » Tue Sep 24, 2013 1:23 am
thanks everyone!
Work hard in Silence, Let Success make the noise.

If you found my Post really helpful, then don't forget to click the Thank/follow me button. :)
Join the discussion